OpenStudy (anonymous):

find dy/dx for the equation xsiny=1

OpenStudy (dumbcow):

use product rule .... also any time you take derivative of a "y" term, multiply by dy/dx \[\rightarrow (1*\sin y)+(x*\cos y \frac{dy}{dx}) = 0\] isolate dy/dx

OpenStudy (anonymous):

thanks!